# A hybrid numerical method for elastic wave propagation in discontinuous media with complex geometry

### v1

@@ Abstract
Keywords: The elastic wave equation, finite difference method, discontinuous Galerkin method, stability, accuracy
@@ 1 Introduction
A method which has seen a lot of applications in the context of high order solvers for wave-dominated problems is the finite difference (FD) method. Finite differences are conceptually easy to understand, computationally efficient and straightforward to implement. However, it was challenging to derive a stable and high order finite difference discretization for initial boundary value problems. This challenge has largely been overcome by using finite difference operators satisfying the summation-by-parts (SBP) property [17]. Operators with this property satisfy a discrete analogue of the integration-by-parts identity, which is the key to derive continuous energy estimates. This method is often successfully combined with the simultaneous-approximation-term (SAT) method to implement boundary conditions via penalization [6], resulting in an SBP-SAT scheme. Alternatively, boundary conditions can be imposed strongly by the projection method [28, 29] or the ghost point method [30]. A drawback of using FD operators however is that they are inherently defined on rectangular grids. To resolve more general geometries, a curvilinear grid is often constructed using a coordinate transformation technique [27, 31]. In certain cases, we need to partition the domain into subdomains that can be more easily mapped to a reference domain, resulting a multi-block finite difference method [24]. The elastic wave equation has been successfully discretized with the multiblock SBP finite difference method, see e.g. [2, 9] for the SAT method and [37, 38] for the ghost point method. The relation between the SAT method and the ghost point method was discussed in [34].
@@
To leverage the computational efficiency of the FD method and the geometric flexibility of the DG method, a hybrid FD-DG method has been proposed. At the FD-DG interface, the discretizations have different induced norms, making it challenging to obtain a stable hybrid method with high order accuracy. This challenge was overcome by Kozdon and Wilcox in [16], where an FD-DG discretization was constructed for the wave equation in first-order form with optimal convergence. It was later shown in the setting of non-conforming FD-FD couplings that in order to obtain optimal convergence when considering problems in second-order form, one needs to use pairs of projection operators [3]. In a recent paper, the work of Kozdon and Wilcox was combined with the notion of projection pairs to solve the wave equation in second-order form, obtaining optimal accuracy [33]. FD-FEM schemes have also been constructed previously for the elastic wave equation [12]. In [21], instead, a first-order velocity-stress formulation was discretized by DG coupled with a staggered grid FD formulation through a finite volume transition zone.
In this work, we follow the coupling methodology in [33] and consider the elastic wave equation in second-order form in a bounded, isotropic and possibly discontinuous material. The first contribution of this paper is an FD-DG discretization of the elastic wave equation in second-order displacement form. We combine an SBP FD discretization with an IPDG discretization using norm compatible projection pairs. It is shown to converge optimally and proven to satisfy a discrete energy estimate under appropriate choice of penalty parameters, guaranteeing stability.
The second contribution is in the optimization of the projection operators for accuracy and spectrum, and its impact on stability. When constructing the projection operators, accuracy conditions are used together with the norm compatibility to set up a system of equations for the coefficients in the projection operators. This system is in general underdetermined and results in a large set of free parameters. We demonstrate that choosing these parameters without optimization, e.g. setting them to zero, yields a stiff system and instead investigate two different strategies of determining them. More specifically, we choose them to minimize the leading order $l^{2}$-error as well as two different measures of spectrum; the Fröbenius norm and an eigenvalue-flattening norm inspired by [16]. This extends the work of [33], where only optimization of accuracy was considered when constructing the operators. The order in which we optimize accuracy and spectrum is also considered. Finally, we demonstrate that the discretization is not sensitive to the distribution of FD and DG nodes along the interface.
@@ 2 The elastic wave equation
where $\mu_{u},\lambda_{u},\mu_{v},\lambda_{v}$ denotes the Lamé parameters of the respective domains, which are assumed to be constant, as well as the respective densities $\rho_{u},\rho_{v}$ which also are constant. For a more detailed description of the above formulation, see e.g. [32]. In order for the problem to be well-posed, we must also impose initial conditions, boundary conditions and suitable interface conditions. To formulate the boundary conditions, we introduce the stress tensors for an arbitrary smooth field $\boldsymbol{u}=[u_{1},u_{2}]^{T}$ as $\sigma(\boldsymbol{u})$
@@
Moreover, let $\boldsymbol{n}$ denote the unit outward normal to $\partial\Omega/\Gamma$. and introduce the tractions $\tau(\boldsymbol{u})=\sigma(\boldsymbol{u})\cdot\boldsymbol{n}$ and $\tau(\boldsymbol{v})=\sigma(\boldsymbol{v})\cdot\boldsymbol{n}$. Then we formulate the traction free boundary conditions as
$\tau(\boldsymbol{u})$ $=\boldsymbol{0}\quad\text{ along }\partial\Omega_{u}/\Gamma,$
$\tau(\boldsymbol{v})$ $=\boldsymbol{0}\quad\text{ along }\partial\Omega_{v}/\Gamma.$
@@
Finally, it should be noted for completeness that the continuous formulation of (1)-(2) can readily be shown to satisfy an energy estimate, see e.g., [5]. We begin in Section 2.1 by introducing the necessary background to discretize (1)-(2) in the SBP-SAT framework, and then in Section 2.2 we discretize it in the IPDG framework.
@@ 2.1 SBP finite difference formulation
Next, we let $D_{1}\approx\partial_{x}$ and $D_{2}\approx\partial_{x}^{2}$ be narrow stencil FD approximations of the first and second derivatives with interior error $\mathcal{O}(h^{2p})$ and $\mathcal{O}(h^{p})$ at a few points along the boundary due to the one-sided stencil. We therefore say that the operator $D_{2}$ is of order $2p$. The SBP property for the differential operators are then defined as follows [23].
@@
Let $D_{1}$ and $D_{2}$ be $2p$’th-order narrow-stencil SBP operators. If $D_{2}$ can be written on the following form
@@
and the remainder $R$ is positive semi-definite, then the operators $D_{1}$ and $D_{2}$ are said to be compatible. Moreover, they are said to be fully compatible if instead
@@
We can then formulate traction operators for constant material parameters in the $x,y$-directions as
@@
The signs of the normals are imposed by multiplying with ±$\pm\mathcal{B}_{x},\pm\mathcal{B}_{y}$ to only pick up the values along the boundaries. This is done as the operators above are defined on the entire mesh $\Omega_{u}$. Using the above notation, we then can formulate the SBP-SAT discretization for (1) as
@@
where $\boldsymbol{u}=[u_{11},...,u_{1n},...,u_{mn}]^{T}$ denotes the grid evaluation of the field with $u_{ij}\approx u(x_{i},y_{j},t)$ and $\alpha$ and $\beta$ are penalty parameters determined by stability analysis.
Note that the last two terms correspond to the SAT-terms which impose the traction-free boundary condition. In the SBP-dG formulation derived in Section 3.2, we only consider the contribution from the last term, more particularly we will only consider the term proportional to $\mathcal{B}_{\Gamma}=I_{2}\otimes I_{x}\otimes B_{0,y}$ as this corresponds to the southern boundary and therefore the interface $\Gamma$. The SBP-SAT discretization (6) is shown to be stable under the choice $\alpha=\beta=-1$ in [10].
@@
where $\mathscr{P}_{q}(K)$ denotes the space of polynomials of order $q$ on the element $K$. Before formulating the weak form of the elastic wave equation, we must introduce a partition of the edges in the triangulation. Following the notation in [8], let $\mathscr{E}_{h}=\{\gamma_{i},i=1,...,n_{e}\}$ denote the set of all edges in $\mathscr{K}_{h}$. Moreover, define an order such that $\gamma_{i}\subset\Omega_{v}$ for $i=1,..,n_{i}$ and $\gamma_{i}\subset\partial\Omega$ for $i=n_{i}+1,..,n_{e}$. Then the definition of interior edges and boundary edges follows naturally as the ordered subsets
@@
To couple the solution between neighbouring elements we shall need the notion of jump and average operators. Let $\gamma\in\mathscr{E}_{h}^{I}$ be an edge shared by the elements $K^{+},K^{-}\in\mathscr{K}_{h}$ with outward pointing normals $\boldsymbol{n}^{+},\boldsymbol{n}^{-}$ respectively and $\varphi$ an arbitrary piecewise smooth function. Then the jump $\llbracket\varphi\rrbracket_{\gamma}$ and average $\{\varphi\}_{\gamma}$ across the edge $\gamma$ are defined depending on if it is an interior or boundary edge of $K^{+}$ accordingly
$\llbracket\varphi\rrbracket_{\gamma}$ $=\varphi_{+}\boldsymbol{n}^{+}+\varphi_{-}\boldsymbol{n}^{-}\qquad\{\varphi\}_{\gamma}$ $=\frac{\varphi_{+}+\varphi_{-}}{2}\qquad\text{ if }\gamma\in\mathscr{E}_{h}^{I}$
$\llbracket\varphi\rrbracket_{\gamma}$ $=\varphi_{+}\boldsymbol{n}^{+}$$\{\varphi\}_{\gamma}=\varphi_{+}\qquad\text{ if }\gamma\in\mathscr{E}_{h}^{\partial}.$
For piecewise smooth vector-valued functions $\boldsymbol{\phi}$, the average is defined in analogous manner as $\{\boldsymbol{\phi}\}_{\gamma}=\frac{\boldsymbol{\phi}^{+}+\boldsymbol{\phi}^{-}}{2}$ for interior edges and $\{\boldsymbol{\phi}\}_{\gamma}=\boldsymbol{\phi}^{+}$ for boundary edges. We do not define a jump operator for $\boldsymbol{\phi}$, if $\boldsymbol{\phi}$ has continuous normal components across the edge $\gamma$ we instead use the jump identity
$\llbracket\varphi\rrbracket_{\gamma}\cdot\{\boldsymbol{\phi}\}_{\gamma}=\varphi^{+}(\boldsymbol{\phi}^{+}\cdot\boldsymbol{n}^{+})+\varphi^{-}(\boldsymbol{\phi}^{-}\cdot\boldsymbol{n}^{-}).$
@@
$\nabla\otimes\nabla=\begin{bmatrix}\partial_{x}^{2}&\partial^{2}_{xy}\\ \partial^{2}_{xy}&\partial_{y}^{2}\end{bmatrix}$
@@
Multiplying with $\boldsymbol{\varphi}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$ and integrating over an element $K\in\mathscr{K}_{h}$ allows us to formulate the problem locally as
@@
$+\sum_{\gamma\in\mathscr{E}_{h}^{I}}\int_{\gamma}\{\tau^{\gamma}(\boldsymbol{v}_{h})\}_{\gamma}\cdot\llbracket\boldsymbol{\varphi}\rrbracket_{\gamma}+\{\tau^{\gamma}(\boldsymbol{\varphi})\}_{\gamma}\cdot\llbracket\boldsymbol{v}_{h}\rrbracket_{\gamma}$
@@
for all $\boldsymbol{\varphi}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$. The first row on the right-hand side is obtained by Green’s identity, together with the first term on the second row, which penalizes the internal traction. The following two terms are added to ensure symmetry and stability of the overall scheme. Notably, they vanish in limit when the mesh size goes to zero, thereby still approximating (2). Finally, the stability parameter $\alpha$ needs to be chosen suitably large for (8) to satisfy an energy estimate [4]. A discussion on the relation between (2) and (8) is presented in [7].
To end this section, we state two estimates needed for the stability analysis of the coupled scheme. First, we bound the terms on the right-hand side of the first row in (8) by the norm of the stress tensor. For notational convenience let $\varepsilon(\boldsymbol{v})$ be the symmetric part of the Jacobian of $\boldsymbol{v}$, called the strain-tensor, and the operation $\sigma(\boldsymbol{v}):\varepsilon(\boldsymbol{v})=\sum_{ij}\sigma_{ij}(\boldsymbol{v})\varepsilon_{ij}(\boldsymbol{v})$ denote matrix contraction, then the following estimate holds.
@@
Let the Lamé parameters $\mu_{v},\lambda_{v}$ satisfy $\lambda_{v}>0$ and $\mu_{v}+\lambda_{v}>0$, then the $L^{2}$ inner-product of the stress tensor $\sigma(\boldsymbol{v})$ and strain tensor $\varepsilon(\boldsymbol{v})$ can be bounded in terms of the $L^{2}$-norm of $\sigma(\boldsymbol{v})$ as follows
$\int_{\Omega_{v}}\sigma(\boldsymbol{v}):\varepsilon(\boldsymbol{v})dA\geq\frac{1}{2\mu_{v}+\lambda_{v}}\|\mathbb{\sigma}(\boldsymbol{v})\|_{L^{2}(\Omega_{v})}^{2}.$ (9)
@@
Using Hooke’s law, we can write the strain tensor in terms of the stress tensor in Voight notation as $\varepsilon_{v}(\boldsymbol{v})=\mathbf{C}\sigma_{v}(\boldsymbol{v})$ where $\mathbf{C}$ is the compliance matrix, $\sigma_{v}(\boldsymbol{v})=[\sigma_{11}(\boldsymbol{v}),\sigma_{22}(\boldsymbol{v}),\sigma_{12}(\boldsymbol{v})]^{T}$ and $\varepsilon_{v}(\boldsymbol{v})=[\varepsilon_{11}(\boldsymbol{v}),\varepsilon_{22}(\boldsymbol{v}),2\varepsilon_{12}(\boldsymbol{v})]^{T}$. In the setting of plane stress, $\mathbf{C}$ is given by
@@
which is clearly symmetric and real from the definition of the Lamé parameters [20]. The desired inequality can therefore be reformulated as finding a lower bound on a constant $k$ such that $\sigma_{v}^{T}\mathbf{C}\sigma_{v}\geq k\sigma_{v}^{T}\sigma_{v}$ holds. From the properties of $\mathbf{C}$ it is clear that $k$ is bounded by the smallest eigenvalue, which can be shown using (10) to yield $k\geq\frac{1}{2(\mu+\lambda)}$. Therefore, it must hold for the choice $k=\frac{1}{2\mu+\lambda}$ from the assumption that $\lambda>0$, and the inequality follows. ∎
@@
Let $\Gamma\subset\partial\Omega$ and introduce $\mathscr{E}_{h}^{\Gamma}\subset\mathscr{E}_{h}^{\partial}$ as the set of edges having non-empty intersection with $\Gamma$. Moreover, let $\mathscr{K}_{h}^{\Gamma}\subset\mathscr{K}_{h}$ denote the set of triangles with at least one edge belonging to $\mathscr{E}_{h}^{\Gamma}$ and $\|\boldsymbol{v}_{h}\|^{2}_{L^{2}(K)}$ the $L^{2}$-norm of the discrete approximation on a triangle $K\in\mathscr{K}_{h}$. Then there exist a positive constant $C^{\prime}>0$ dependent on the quality of $\mathscr{K}_{h}$ such that
$\|\sigma(\boldsymbol{v}_{h})\cdot\boldsymbol{n}\|^{2}_{L^{2}(\Gamma)}$ $\leq\sum_{\gamma\in\mathscr{E}_{h}^{\Gamma}}\|\sigma(\boldsymbol{v}_{h})\cdot\boldsymbol{n}\|^{2}_{L^{2}(\gamma)}\leq\sum_{K\in\mathscr{K}_{h}^{\Gamma}}\|\sigma(\boldsymbol{v}_{h})\cdot\boldsymbol{n}\|^{2}_{L^{2}(\partial K)}$

### v2

@@ Abstract
Keywords: finite difference methods, discontinuous Galerkin methods, hybrid methods , elastic wave equation, high order methods
MSC: 65M06, 65M60, 65M12
@@ 1 Introduction
A method which has seen a lot of applications in the context of high order solvers for wave-dominated problems is the finite difference (FD) method. Finite differences are conceptually easy to understand, computationally efficient and straightforward to implement. However, it was challenging to derive a stable and high order finite difference discretization for initial boundary value problems. This challenge has largely been overcome by using finite difference operators satisfying the summation-by-parts (SBP) property [19]. Operators with this property satisfy a discrete analogue of the integration-by-parts identity, which is the key to derive continuous energy estimates. This method is often successfully combined with the simultaneous-approximation-term (SAT) method to implement boundary conditions via penalization [7], resulting in an SBP-SAT scheme. Alternatively, boundary conditions can be imposed strongly by the projection method [30, 31] or the ghost point method [32]. A drawback of using FD operators however is that they are inherently defined on rectangular grids. To resolve more general geometries, a curvilinear grid is often constructed using a coordinate transformation technique [29, 33]. In certain cases, we need to partition the domain into subdomains that can be more easily mapped to a reference domain, resulting in a multi-block finite difference method [26]. The elastic wave equation has been successfully discretized with the multiblock SBP finite difference method, see e.g. [2, 10] for the SAT method and [39, 40] for the ghost point method. The relation between the SAT method and the ghost point method was discussed in [36]. It should be noted that the aforementioned papers considers the equation in second order form as this is known to, in general, result in fewer unknowns than first order formulations and improved order of convergence for SBP-SAT discretizations.
@@
To leverage the computational efficiency of the FD method and the geometric flexibility of the DG method, a hybrid FD-DG method has been proposed. At the FD-DG interface, the discretizations have different induced norms, making it challenging to obtain a stable hybrid method with high order accuracy. This challenge was overcome by Kozdon and Wilcox in [18], where an FD-DG discretization was constructed for the wave equation in first-order form with optimal convergence. It was later shown in the setting of non-conforming FD-FD couplings that in order to obtain optimal convergence when considering problems in second-order form, one needs to use pairs of projection operators [3]. In a recent paper, the work of Kozdon and Wilcox was combined with the notion of projection pairs to solve the wave equation in second-order form, obtaining optimal accuracy [35]. FD-FEM schemes have also been constructed previously for Maxwell’s equation [6] and the elastic wave equation [13]. In [23], instead, a first-order velocity-stress formulation was discretized by DG coupled with a staggered grid FD formulation through a finite volume transition zone.
In this work, we follow the coupling methodology in [35] and consider the elastic wave equation in second-order form in a bounded, isotropic and possibly discontinuous material. The first contribution of this paper is an FD-DG discretization of the elastic wave equation in second-order displacement form. We combine an SBP FD discretization with an interior penalty DG (IPDG) discretization using norm compatible projection pairs. It is shown to converge optimally and proven to satisfy a discrete energy estimate under appropriate choice of penalty parameters, guaranteeing stability.
The second contribution is in the optimization of the projection operators for accuracy and spectrum, and its impact on stability. When constructing the projection operators, accuracy conditions are used together with the norm compatibility to set up a system of equations for the coefficients in the projection operators. This system is in general underdetermined and results in a large set of free parameters. We demonstrate that choosing these parameters without optimization, e.g. setting them to zero, yields a stiff system and instead investigate two different strategies of determining them. More specifically, we choose them to minimize the leading order $l^{2}$-error as well as two different measures of spectrum; the Frobenius norm and an eigenvalue-flattening norm inspired by [18]. This extends the work of [35], where only optimization of accuracy was considered when constructing the operators. The order in which we optimize accuracy and spectrum is also considered. Finally, we demonstrate that the discretization is not sensitive to the distribution of FD and DG nodes along the interface.
@@ 2 The elastic wave equation
where $\mu_{u},\lambda_{u},\mu_{v},\lambda_{v}$ denotes the Lamé parameters of the respective domains, which are assumed to be constant, as well as the respective densities $\rho_{u},\rho_{v}$ which also are constant. For a more detailed description of the above formulation, see e.g. [34]. In order for the problem to be well-posed, we must also impose initial conditions, boundary conditions and suitable interface conditions. To formulate the boundary conditions, we introduce the stress tensor for an arbitrary smooth field $\boldsymbol{u}=[u_{1},u_{2}]^{T}$ as $\sigma(\boldsymbol{u})$
@@
Moreover, let $\boldsymbol{n}$ denote the unit outward normal to $\partial\Omega/\Gamma$. and introduce the tractions $\boldsymbol{\tau}(\boldsymbol{u})=\sigma(\boldsymbol{u})\cdot\boldsymbol{n}$ and $\boldsymbol{\tau}(\boldsymbol{v})=\sigma(\boldsymbol{v})\cdot\boldsymbol{n}$. Then we formulate the traction free boundary conditions as
$\boldsymbol{\tau}(\boldsymbol{u})$ $=\boldsymbol{0}\quad\text{ along }\partial\Omega_{u}/\Gamma,$
$\boldsymbol{\tau}(\boldsymbol{v})$ $=\boldsymbol{0}\quad\text{ along }\partial\Omega_{v}/\Gamma.$
@@
Finally, it should be noted for completeness that the continuous formulation of (1)-(2) can readily be shown to satisfy an energy estimate, see e.g. [5]. We begin in Section 2.1 by introducing the necessary background to discretize (1)-(2) in the SBP-SAT framework, and then in Section 2.2 we discretize it in the IPDG framework.
@@ 2.1 SBP finite difference formulation
Next, we let $D_{1}\approx\partial_{x}$ and $D_{2}\approx\partial_{x}^{2}$ be narrow stencil FD approximations of the first and second derivatives with interior error $\mathcal{O}(h^{2p})$ and $\mathcal{O}(h^{p})$ at a few points along the boundary due to the one-sided stencil. We then say that the operator $D_{2}$ is of order $2p$. The SBP property for the differential operators are then defined as follows [25].
@@
Let $D_{1}$ and $D_{2}$ be $2p$’th-order narrow-stencil SBP operators. If $D_{2}$ can be written in the following form
@@
and the remainder $R$ is symmetric positive semi-definite, then the operators $D_{1}$ and $D_{2}$ are said to be compatible. Moreover, they are said to be fully compatible if instead
@@
To formulate the traction, we introduce the following operators
@@
so that the tractions on the boundaries are obtained by multiplying with ±$\pm\mathcal{B}_{x},\pm\mathcal{B}_{y}$ to only pick up the values along the boundaries. This is done as the operators are defined on the entire mesh $\Omega_{u}$. Using the above notation, we then can formulate the SBP-SAT discretization for (1) as
@@
where $\boldsymbol{u}=[u_{11},...,u_{1n},...,u_{mn}]^{T}$ denotes the grid evaluation of the field with $u_{ij}\approx u(x_{i},y_{j},t)$ and $\alpha,\beta$ are penalty parameters determined by stability analysis.
Note that the last two terms correspond to the SAT-terms which impose the traction-free boundary condition. In the SBP-DG formulation derived in Section 3.2, we only consider the contribution from the term that corresponds to the southern boundary and therefore the interface $\Gamma$. The SBP-SAT discretization (6) is shown to be stable under the choice $\alpha=\beta=-1$ in [11].
@@ 2.2 Discontinuous Galerkin formulation
where $\mathscr{P}_{q}(K)$ denotes the space of polynomials of order $q$ on the element $K$. Before formulating the IPDG form of the elastic wave equation, we must introduce a partition of the edges in the triangulation. Following the notation in [9], let $\mathscr{E}_{h}=\{\gamma_{i},i=1,...,n_{e}\}$ denote the set of all edges in $\mathscr{K}_{h}$. Moreover, define an order such that $\gamma_{i}\subset\Omega_{v}$ for $i=1,..,n_{i}$ and $\gamma_{i}\subset\partial\Omega$ for $i=n_{i}+1,..,n_{e}$. Then the definition of interior edges and boundary edges follows naturally as the ordered subsets of $\mathscr{E}_{h}$
@@
To couple the solution between neighbouring elements we shall need the notion of jump and average operators. Let $\gamma\in\mathscr{E}_{h}^{I}$ be an edge shared by the elements $K^{+},K^{-}\in\mathscr{K}_{h}$ with unit outward pointing normals $\boldsymbol{n}^{+},\boldsymbol{n}^{-}$ respectively and $\varphi$ an arbitrary piecewise smooth function. Then the jump $\llbracket\varphi\rrbracket_{\gamma}$ and average $\{\varphi\}_{\gamma}$ across the edge $\gamma$ are defined depending on if it is an interior or boundary edge of $K^{+}$ accordingly
$\llbracket\boldsymbol{\varphi}\rrbracket_{\gamma}$ $=\varphi_{+}\boldsymbol{n}^{+}+\varphi_{-}\boldsymbol{n}^{-},\qquad\{\varphi\}_{\gamma}$ $=\frac{\varphi_{+}+\varphi_{-}}{2}\qquad\text{ if }\gamma\in\mathscr{E}_{h}^{I},$
$\llbracket\boldsymbol{\varphi}\rrbracket_{\gamma}$ $=\varphi_{+}\boldsymbol{n}^{+},$$\{\varphi\}_{\gamma}=\varphi_{+},\qquad\text{ if }\gamma\in\mathscr{E}_{h}^{\partial}.$
For piecewise smooth vector-valued functions $\boldsymbol{\phi}$, the average is defined in analogous manner as $\{\boldsymbol{\phi}\}_{\gamma}=\frac{\boldsymbol{\phi}^{+}+\boldsymbol{\phi}^{-}}{2}$ for interior edges and $\{\boldsymbol{\phi}\}_{\gamma}=\boldsymbol{\phi}^{+}$ for boundary edges. Likewise, we set the jump to be $\llbracket\boldsymbol{\phi}\rrbracket_{\gamma}=\frac{\boldsymbol{\phi}^{+}-\boldsymbol{\phi}^{-}}{2}$ for interior edges and $\llbracket\boldsymbol{\phi}\rrbracket_{\gamma}=\boldsymbol{\phi}^{+}$ for boundary edges.
@@
$\nabla\otimes\nabla=\begin{bmatrix}\partial_{x}^{2}&\partial^{2}_{xy}\\ \partial^{2}_{xy}&\partial_{y}^{2},\end{bmatrix}$
@@
Multiplying with $\boldsymbol{\varphi}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$ and integrating over an element $K\in\mathscr{K}_{h}$ allows us to formulate the problem locally on $K$ as
@@
$+\sum_{\gamma\in\mathscr{E}_{h}^{I}}\int_{\gamma}\{\boldsymbol{\tau}^{\gamma}(\boldsymbol{v}_{h})\}_{\gamma}\cdot\llbracket\boldsymbol{\varphi}\rrbracket_{\gamma}+\{\boldsymbol{\tau}^{\gamma}(\boldsymbol{\varphi})\}_{\gamma}\cdot\llbracket\boldsymbol{v}_{h}\rrbracket_{\gamma}$
@@
for all $\boldsymbol{\varphi}\in\mathscr{V}_{h}^{q}\times\mathscr{V}_{h}^{q}$. Here the operation $A:B=\sum_{ij}A_{ij}B_{ij}$ denotes matrix contraction. The first row on the right-hand side is obtained by Green’s identity, together with the first term on the second row. The following two terms are added to ensure symmetry and stability of the overall scheme. Notably, they vanish in limit when the mesh size goes to zero, thereby still approximating (2). Finally, the stability parameter $\alpha$ needs to be chosen suitably large for (8) to satisfy an energy estimate [4]. A discussion on the relation between (2) and (8) is presented in [8].
To end this section, we state two estimates needed for the stability analysis of the coupled scheme. First, we bound the terms on the right-hand side of the first row in (8) by the norm of the stress tensor. For notational convenience let $\varepsilon(\boldsymbol{v})$ be the symmetric part of the Jacobian of $\boldsymbol{v}$, called the strain-tensor, then the following estimate holds.
@@
Let the Lamé parameters $\mu_{v},\lambda_{v}$ be positive and consider the case of plane strain, then the $L^{2}$ inner-product of the stress tensor $\sigma(\boldsymbol{v})$ and strain tensor $\varepsilon(\boldsymbol{v})$ can be bounded in terms of the $L^{2}$-norm of $\sigma(\boldsymbol{v})$ as follows
$\int_{\Omega_{v}}\sigma(\boldsymbol{v}):\varepsilon(\boldsymbol{v})dA\geq\frac{1}{4\mu_{v}+2\lambda_{v}}\|\mathbb{\sigma}(\boldsymbol{v})\|_{L^{2}(\Omega_{v})}^{2},$ (9)
@@
Using Hooke’s law, we can write the strain tensor in terms of the stress tensor in Voigt notation as $\boldsymbol{\varepsilon}_{v}(\boldsymbol{v})=\mathbf{C}\boldsymbol{\sigma}_{v}(\boldsymbol{v})$ where $\mathbf{C}$ is the compliance matrix, $\boldsymbol{\sigma}_{v}(\boldsymbol{v})=[\sigma_{11}(\boldsymbol{v}),\sigma_{22}(\boldsymbol{v}),\sigma_{12}(\boldsymbol{v})]^{T}$ and $\boldsymbol{\varepsilon}_{v}(\boldsymbol{v})=[\varepsilon_{11}(\boldsymbol{v}),\varepsilon_{22}(\boldsymbol{v}),2\varepsilon_{12}(\boldsymbol{v})]^{T}$. In the setting of plane strain, $\mathbf{C}$ is given by
@@
which is clearly symmetric and real from the definition of the Lamé parameters [22]. The desired inequality can therefore be reformulated as finding a lower bound on a constant $k$ such that $\boldsymbol{\sigma}_{v}^{T}\mathbf{C}\boldsymbol{\sigma}_{v}\geq k\boldsymbol{\sigma}_{v}^{T}\boldsymbol{\sigma}_{v}$ holds. From the properties of $\mathbf{C}$ it is clear that $k$ is bounded by the smallest eigenvalue, which can be shown using (10) to yield $k\leq\frac{1}{2(\mu+\lambda)}$. Therefore, it must in particular hold for the choice $k=\frac{1}{4\mu+2\lambda}$ by the assumption on positivity of the Lamé parameters. ∎
@@
Let $\Gamma\subset\partial\Omega$ and introduce $\mathscr{E}_{h}^{\Gamma}\subset\mathscr{E}_{h}^{\partial}$ as the set of boundary edges having non-empty intersection with $\Gamma$. Moreover, let $\mathscr{K}_{h}^{\Gamma}\subset\mathscr{K}_{h}$ denote the set of triangles with at least one edge belonging to $\mathscr{E}_{h}^{\Gamma}$ and $\|\boldsymbol{v}_{h}\|^{2}_{L^{2}(K)}$ the $L^{2}$-norm of the discrete approximation on a triangle $K\in\mathscr{K}_{h}$. Then there exist a positive constant $C^{\prime}>0$ dependent on the quality of $\mathscr{K}_{h}$ such that
$\|\sigma(\boldsymbol{v}_{h})\cdot\boldsymbol{n}\|^{2}_{L^{2}(\Gamma)}$ $=\sum_{\gamma\in\mathscr{E}_{h}^{\Gamma}}\|\sigma(\boldsymbol{v}_{h})\cdot\boldsymbol{n}\|^{2}_{L^{2}(\gamma)}\leq\sum_{K\in\mathscr{K}_{h}^{\Gamma}}\|\sigma(\boldsymbol{v}_{h})\cdot\boldsymbol{n}\|^{2}_{L^{2}(\partial K)}$